# Second- and third-order properties of multidimensional Langevin equations

### 25 Dec 2023 v5

@@ 7.4. Nonlinear drift
(253) ${{}\widehat{\mathbf{A}}^{n}}^{y(i)}_{y(j)}={{}\widehat{\mathbf{A}}^{n}}^{(i)}_{(j)}.$
We also have, for $m\geq 3$:
(254) $\left|\Delta y(\tau)\mathbf{x}^{(m)}(\tau)\mathbf{x}(0)^{\mathsf{T}}\right|$ $\leq\mathcal{O}(a,b)$
which implies, for $n\geq 1$ ($m\geq 3$):
(256) $\left|{{}\widehat{\mathbf{A}}^{n}}^{y(m)}\right|\leq\mathcal{O}(a,b).$
@@
(259) ${{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{\pm}\langle\mathbf{x}^{\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle+{{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{(2)\pm}\langle\mathbf{x}^{(2)\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle$ $=-\langle\mathbf{x}^{\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{\pm}}^{\mathsf{T}}-\langle\mathbf{x}^{\pm}{\mathbf{x}^{(2)\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{(2)\pm}}^{\mathsf{T}}+a\mathcal{O}((a,b)^{2}),$
(260) ${{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{\mp}\langle\mathbf{x}^{\mp}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle+{{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{(2)\mp}\langle\mathbf{x}^{(2)\mp}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle$ $=\langle\mathbf{x}^{\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)\mp}_{\pm}}^{\mathsf{T}}+\langle\mathbf{x}^{\pm}{\mathbf{x}^{(2)\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)\mp}_{(2)\pm}}^{\mathsf{T}}+a\mathcal{O}((a,b)^{2}).$
This shows the time-reversibility of the covariance function $\langle\Delta y(\tau)\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$. For $\langle\Delta y(\tau)\mathbf{x}^{(2)}(0)\rangle$, we need to show that for $n\geq 1$:
@@
Here, we need to use:
@@
(263) $\left|{{}\widehat{\mathbf{A}}^{n}}^{y(m)}_{(0)}\right|$ $\leq\mathcal{O}(a,b),\quad m\geq 4,$
(264) $\left|{{}\widehat{\mathbf{A}}^{n}}^{y(2)}_{(m)}\right|$ $\leq\mathcal{O}(a,b),\quad m\geq 4.$
@@
(265) ${{}\widehat{\mathbf{A}}}^{(3)}_{(0)}+{{}\widehat{\mathbf{A}}}^{(3)}_{(3)}\langle\mathbf{x}^{(3)}\rangle=\mathbf{0}.$

### 11 Jan 2024 v6

@@ 7.5. Integrated variables
(257) ${{}\widehat{\mathbf{A}}^{n}}^{y(i)}_{y(j)}={{}\widehat{\mathbf{A}}^{n}}^{(i)}_{(j)},$
and for $i\geq 1$:
(258) ${{}\widehat{\mathbf{A}}^{n}}^{y(i)}_{(i)}=\mathcal{O}(a,b).$
We also have:
(259) ${{}\widehat{\mathbf{A}}^{n}}^{(1)}_{(0)}$ $=a^{2}\mathcal{O}(a,b),$
(260) ${{}\widehat{\mathbf{A}}^{n}}^{y(0)}_{(0)}$ $=a^{2}\mathcal{O}(a,b).$
From orthogonality of Hermite polynomials and symmetry, we have:
(261) $\langle\Delta y(\tau)\mathbf{x}^{(2)}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ $=\mathcal{O}((a,b)^{2}),$
(262) $\langle\Delta y(\tau)\mathbf{x}^{(3)}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ $=\mathcal{O}(a,b),$
which implies:
(263) ${{}\widehat{\mathbf{A}}^{n}}^{y(2)}$ $=\mathcal{O}((a,b)^{2}),$
(264) ${{}\widehat{\mathbf{A}}^{n}}^{y(3)}$ $=\mathcal{O}(a,b).$
@@
(267) ${{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{\pm}\langle\mathbf{x}^{\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle+{{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{(2)\pm}\langle\mathbf{x}^{(2)\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle$ $=-\langle\mathbf{x}^{\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{\pm}}^{\mathsf{T}}-\langle\mathbf{x}^{\pm}{\mathbf{x}^{(2)\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{(2)\pm}}^{\mathsf{T}}+a\mathcal{O}((a,b)^{2})$
(268) ${{}\widehat{\mathbf{A}}^{n}}^{y(1)+}_{-}\langle\mathbf{x}^{-}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle+{{}\widehat{\mathbf{A}}^{n}}^{y(1)+}_{(2)-}\langle\mathbf{x}^{(2)-}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle$ $=\langle\mathbf{x}^{+}{\mathbf{x}^{+}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)-}_{+}}^{\mathsf{T}}+\langle\mathbf{x}^{+}{\mathbf{x}^{(2)+}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)-}_{(2)+}}^{\mathsf{T}}+a\mathcal{O}((a,b)^{2}),$
where the base case $n=1$ for Eq. (267) follows from $L(\mathbf{x}^{(2)+},y)=0$ and for Eq. (268) follows from Eqs. (234) and (240). This shows the time-reversibility of the covariance function $\langle\Delta y(\tau)\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$. For $\langle\Delta y(\tau)\mathbf{x}^{(2)}(0)\rangle$, we need to show that for $n\geq 1$:
@@
where the base case $n=1$ for the upper sign follows from $L(\mathbf{x}^{(2)+},y)=0$ and for the lower sign follows from Eqs. (234) and (240). Here, we need to use:
(270) ${{}\widehat{\mathbf{A}}^{n}}^{(2)}_{(0)}$ $=a\mathcal{O}(a,b),$
(271) $\left|{{}\widehat{\mathbf{A}}^{n}}^{(m)}_{(0)}\right|$ $\leq\mathcal{O}((a,b)^{2}),\quad m\geq 4$
@@
(273) ${{}\widehat{\mathbf{A}}^{n}}^{y(4)}_{(0)}$ $=\mathcal{O}(a,b).$
@@
(274) ${{}\widehat{\mathbf{A}}}^{(3)}_{(0)}+{{}\widehat{\mathbf{A}}}^{(3)}_{(3)}\langle\mathbf{x}^{(3)}\rangle=a\mathcal{O}((a,b)^{2}).$
@@
7.6. Second-order covariance functions to quadratic order
We now consider the case where second-order time-antisymmetric quantities are $a^{2}\mathcal{O}((a,b)^{2})$ and third-order time-antisymmetric quantities are $a\mathcal{O}((a,b)^{2})$. We seek to show that second-order covariance functions obey time-reversal symmetry, with $a^{2}\mathcal{O}((a,b)^{2})$ corrections. We need to show that, for $n\geq 1$:
(275) ${{}\widehat{\mathbf{A}}^{n}}^{\pm}_{\pm}\langle\mathbf{x}^{\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle+{{}\widehat{\mathbf{A}}^{n}}^{\pm}_{(2)\pm}\langle\mathbf{x}^{(2)\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle$ $=\langle\mathbf{x}^{\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{\pm}_{\pm}}^{\mathsf{T}}+\langle\mathbf{x}^{\pm}{\mathbf{x}^{(2)\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{\pm}_{(2)\pm}}^{\mathsf{T}}+a^{2}\mathcal{O}((a,b)^{2}),$
(276) ${{}\widehat{\mathbf{A}}^{n}}^{+}_{-}\langle\mathbf{x}^{-}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle+{{}\widehat{\mathbf{A}}^{n}}^{+}_{(2)-}\langle\mathbf{x}^{(2)-}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle$ $=-\langle\mathbf{x}^{+}{\mathbf{x}^{+}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{-}_{+}}^{\mathsf{T}}-\langle\mathbf{x}^{+}{\mathbf{x}^{(2)+}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{-}_{(2)+}}^{\mathsf{T}}+a^{2}\mathcal{O}((a,b)^{2}),$
where the base case $n=1$ is assumed for Eq. (275), and for Eq. (276) follows from our assumptions that $\mathbf{D}^{+-}$ and $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle$ are $a^{2}\mathcal{O}((a,b)^{2})$, and that $\langle\mathbf{x}^{(2)\pm}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle=a\mathcal{O}((a,b)^{2})$. Using the properties stated in the subsection “Nonlinear drift”, the induction step is readily performed. For the case of an integrated variable $y$ of even parity, we additionally need the equalities in the subsection “Integrated variables” and:
(277) ${{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{(0)}+{{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{(2)+}\langle\mathbf{x}^{(2)+}\rangle+{{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{(3)+}\langle\mathbf{x}^{(3)+}\rangle=\mp\langle\mathbf{x}^{\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y}_{\pm}}^{\mathsf{T}}\mp\langle\mathbf{x}^{\pm}{\mathbf{x}^{(2)\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y}_{(2)\pm}}^{\mathsf{T}}+a^{2}\mathcal{O}((a,b)^{2}).$
We also need to use Eq. (274) and:
(278) ${{}\widehat{\mathbf{A}}}^{(2)}_{(2)}\langle\mathbf{x}^{(2)}\rangle+{{}\widehat{\mathbf{A}}}^{(2)}_{(3)}\langle\mathbf{x}^{(3)}\rangle=a^{2}\mathcal{O}((a,b)^{2}).$